OpenStudy (anonymous):

Write an equation in slope-intercept form for the the line that is parallel to 3x+2y=12 and contains the point (6,-2).

OpenStudy (confluxepic):

\[\huge \color{blue}{Welcome~To~OpenStudy!!!}\]

OpenStudy (confluxepic):

Since the lines have to be parallel the slopes also have to be the same.

OpenStudy (anonymous):

The slope-intercept form is an equation in this form:\(\ y = mx + b\), where\(\ m\) is the slope and\(\ b\) is the location where the line touches the y-axis when\(\ x=0\). In a line that is parallel to\(\ y = mx + b\) the only thing that changes is\(\ b\). It means that the slope\(\ m\) is the same but it touches the y-axis in a different place when\(\ x=0\).

OpenStudy (anonymous):

I know, I got the answer an y=-3/2x+6 but instead of 6 it should be 7 but I want to know how

OpenStudy (anonymous):

The equation that was given cannot me further manipulated. You must end up with 2 equation: that that was given (written now in slope-intercept form) and the new one that you found.

OpenStudy (confluxepic):

True.

OpenStudy (anonymous):

So I cannot go farther than the equation?